3 moles of a gas mixture having volume v and temperature t is compressed to 1/5th of the initial volume.FInd the change in adiabatic compressibility if the gas obeys PV^19/13 constant. (R=8.3J).

Grade:12

1 Answers

Harishwar IIT Roorkee
9 years ago
Adiabatic compressibility is defined: [\beta_S=-\frac{1}{V}\left(\frac{\partial V}{\partial p}\right)_S]